OpenStudy (solomonzelman):

Log(x)=Log(2) Idk please help!

OpenStudy (lukecrayonz):

Oh Soloman :-) x=2!

OpenStudy (solomonzelman):

Why OH?

OpenStudy (solomonzelman):

Anf why 2?

OpenStudy (shamil98):

log(x) = log(2) To get x isolated divide both sides by the log . \[\large \frac{ log(x) }{ \log } = \frac{ \log(2) }{ \log }\] they cancel \[\large x = 2\]

OpenStudy (solomonzelman):

Nice pic, shamil.

OpenStudy (solomonzelman):

Well, i was kidding, and putting it like this is not really valid, you don't canlce logs

OpenStudy (solomonzelman):

\[\log(a)=\log(b)~~~~~\therefore~~~a=b\]

OpenStudy (solomonzelman):

identity.

OpenStudy (skullpatrol):

Obviously, 2 is the only number that will make it a true statement.

OpenStudy (solomonzelman):

Tnx every one came to my question, preciate your help.

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

I am bored.

OpenStudy (solomonzelman):

I'll just draw, don't pay attention....

OpenStudy (solomonzelman):